A ball of mass 8kg falls from rest from a height of 100m. Neglecting air resistance, calculate its kinetic energy after falling a distance of 30m .(Take g as 10ms^-2)


1
Expert's answer
2021-02-17T10:41:39-0500
"K=mg(H-h)\\K=(8)(9.8)(100-30)=5488\ J"